Dividends Per Share

Lightfoot Inc., a software development firm, has Shares of ownership of a corporation.stock outstanding as follows: 10,000 shares of cumulative A class of stock with preferential rights over common stock.preferred 4% stock, $20 par, and 13,000 shares of $100 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $3,000; second year, $5,000; third year, $27,830; fourth year, $48,170.

Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $3000 (WRONG) $5000 (WRONG) $16000(WRONG) $
Common stock (dividend per share) $ $ $ $
Answer #1

Working:

For 1st & 2nd year the dividend on preferred stock cannot be given in full ss it is less than what is being distributed.

Proportionate dividend us allocated to both shares in 1st and 2nd year.

Preferred stock value= 10000 x $ 20 = $ 200000

Common stock value = 13000 x $ 100 = $ 1300000

Proportion of preferred stock to common stock = $200000/$1500000 x 100

= 13.33%

Common stock share = 86.67%

For third year and fourth year where dividend distributed is more than preferred dividend, entire preferred dividend can be given. Preferred dividend must be given first and then common stock dividend.

1st year 2nd year 3rd year 4th year
Preferred stock (dividend per shate) (3000 x 13.33%)/10000 = 0.04 (5000 x 13.33%)/10000 = $ 0.067 $200000 x 4% = $ 8000 $ 200000 x 4% = $ 8000
Common stock (dividend per share) ($ 3000 - $ 400)/13000 = $ 0.20 ($ 5000 - $667)/13000 = $ 0.33 $27830 - $ 8000 = $ 19830 $ 48170 - $ 8000 = $ 40170